OpenStudy (anonymous):

write each ratio as a fraction and simplify 1/4ft to 3/2ft

OpenStudy (anonymous):

1/4 = 3/2 = 2/12 = 1/6?

OpenStudy (anonymous):

Can someone check this please?

OpenStudy (anonymous):

do you mean 1/4 as 1:4

OpenStudy (anonymous):

no thats how it looks

OpenStudy (amistre64):

1/4 1/4 ---- = ----- = 1/6 3/2 6/4

OpenStudy (anonymous):

Thank you!

OpenStudy (amistre64):

the 4s cancel casue they are the same top to bottm and the 1/6 is left :) little simpler than trying to reduce and such

OpenStudy (anonymous):

ha well what I did matches yours lol